# Has an AI model been trained on your images?

### v1

@@ 4 Discussion
We hypothesize a few different mechanisms that may explain why generative-AI models behave this way. One possibility is that when a seed image is partially corrupted with additive noise (proportional to the user-supplied strength parameter) and placed in the latent space for denoising, because of previous exposure to a training image-caption pair, the previously learned local gradients guide the denoising to a latent representation near a trained image. This explanation would be consistent with the differences seen in Fig. 2(a)-(d), where a memorized image/caption yields more self-similar images than an in-training image/caption, which yields more self-similar images than an in-training image/alternate caption. That is, the level of exposure to a specific image/caption pair at training leads to proportionally learned gradients in the denoiser.
@@
Understanding why models are biased to produce content similar to their training data may provide insights into reducing the likelihood of direct copyright infringement in the form of reproducing training data, and may provide insights into how a model can be made to forget training exemplars.
An attractive aspect of our membership inference for generative image models is that it does not require direct access to the details of the model architecture or trained weights, is computationally efficient, and generalizes to multiple different AI models. A drawback of our approach is that it only applies to models that allow for an image-to-image synthesis with a controllable strength parameter, as compared to text-to-image. Depending on the underlying mechanism by which models produce images similar to their in-training data, our method may be adaptable to text-to-image generation.
@@
Data Availability
The STROLL, Carlini, and Midjourney datasets are available at: http://doi.org/10.5281/zenodo.14573149.
@@ Acknowledgments
This work was supported by a grant from the University of California Noyce Initiative. Thanks to Nicholas Carlini and Emily Cooper for their insightful comments and suggestions.

### v2

@@ 4 Discussion
We hypothesize a few different mechanisms that may explain why generative-AI models behave this way. One possibility is that when a seed image is partially corrupted with additive noise (proportional to the user-supplied strength parameter) and placed in the latent space for denoising, because of previous exposure to a training image-caption pair, the previously learned local gradients guide the denoising to a latent representation near a trained image. This explanation would be consistent with the differences seen in Figure 3(a)-(d), where a memorized image/caption yields more self-similar images than an in-training image/caption, which yields more self-similar images than an in-training image/alternate caption. That is, the level of exposure to a specific image/caption pair at training leads to proportionally learned gradients in the denoiser.
@@
Understanding why models are biased to produce content similar to their training data may provide insights into reducing the likelihood of infringement in the form of reproducing training data, and may provide insights into how a model can be made to forget training exemplars.
An attractive aspect of our membership inference for generative image models is that it does not require access to model architecture details or trained weights, is computationally efficient, and generalizes to multiple different AI models. A drawback of our approach is that it only applies to models that allow for an image-to-image synthesis with a controllable strength parameter, as compared to text-to-image. Depending on the underlying mechanism by which models produce images similar to their in-training data, our method may be adaptable to text-to-image generation.
@@ Acknowledgments
This work was funded by funding from the University of California Noyce Initiative. The authors thank Nicholas Carlini for providing the list of extracted instances of memorization from [3], and both Nicholas Carlini and Milad Nasr for helpful discussions.